# Identifying heterogeneous micromechanical properties of biological tissues via physics-informed neural networks

### v1

@@ 5.1 Reference data generation
Herein, $\lambda$ and $\mu$ are the Lamé’s elasticity parameters. Further, $\Psi$ is the strain energy density function expressed as
where $I_{1}=\text{trace}(F^{T}\cdot F)$, $F_{ij}=\delta_{ij}+u_{i,j}$, and $u_{i,j}$ is the displacement vector. The Green-Lagrangian strain is defined as
$\epsilon_{ij}=\frac{1}{2}(F_{ki}F_{kj}-\delta_{ij}).$
In the current work, we followed a plane-strain formulation:
$\lambda=\frac{E\nu}{(1+\nu)(1-2\nu)},\quad\textrm{and}\quad\mu=\frac{E}{2(1+\nu)},$
where we assumed $\nu=0.3$. The finite element mesh was deformed by applying an equibiaxial displacement ($d=0.2$) to the boundaries, as shown in Fig. 1.
@@ 5.3 PINNs for heterogeneous material properties identification
As an extension to prior work on utilizing PINNs to determine the macroscopic constitutive parameters in linear and hyperelastic materials [54, 55], we demonstrate a proof-of-concept study to identify the full-field elastic properties across the 2D geometric domain. Our work aims to facilitate future studies to better understand the interconnection between material microscopic properties and its macroscopic mechanical functions. All PINNs models were constructed using the DeepXDE library [85], and the code is publicly available in the GitHub repository https://github.com/lu-group/pinn-heterogeneous-material.
@@ 5.3.2 Loss functions
Author contributions
W. W.: Conceptualization, Methodology, Formal analysis, Visualization, Writing – Original Draft; M. D.: Conceptualization, Writing – Review & Editing; K. T. T.: Conceptualization, Writing – Review & Editing; M. A. J.: Conceptualization, Resources, Writing – Review & Editing; L. L.: Conceptualization, Writing – Review & Editing, Supervision. All authors discussed the results.
@@ Acknowledgments
This work was supported by the Cora Topolewksi Pediatric Valve Center at the Children’s Hospital of Philadelphia, an Additional Ventures Expansion Award, the National Institutes of Health, and the U.S. Department of Energy. WW was initially supported by NHLBI T32 HL007915 and transitioned to NHLBI K25 HL168235. MAJ was supported by NIH R01 HL153166, an Additional Ventures Single Ventricle Grant, and the Topolewski Endowed Chair in Pediatric Cardiology. MD and LL were supported by the U.S. Department of Energy [DE-SC0022953].
@@
The code for this study is available in the GitHub repository https://github.com/lu-group/pinn-heterogeneous-material.

### v3

@@ 5.2 Hyperelastic constitutive models
The plane-strain formulation of the Neo-Hookean model is expressed as
$\lambda=\frac{E\nu}{(1+\nu)(1-2\nu)},\quad\textrm{and}\quad\mu=\frac{E}{2(1+\nu)}.$
In the case of plane stress, the Lamé constant $\lambda$ becomes
$\lambda=\frac{2\lambda\mu}{\lambda+2\mu}.$
Mooney Rivlin model. The Mooney Rivlin model [81] is made of a linear combination of the first and second principal invariants, $I_{1}$ and $I_{2}$. The strain energy density function is expressed as
$\Psi(I_{1},I_{2})=\frac{\mu_{1}}{2}(I_{1}-2)+\frac{\mu_{2}}{2}(I_{2}-2),$
where $\mu_{1}$ and $\mu_{2}$ are empirically determined material constants. The second principal invariant is defined as $I_{2}=\frac{1}{2}[I_{1}^{2}-\text{trace}(C\cdot C)]$, where $C$ is the right Cauchy-Green tensor defined as $C=F^{T}\cdot F$. Typically, $\mu_{2}$ is much smaller than $\mu_{1}$. Further, for small deformation problems, the sum of $\mu_{1}$ and $\mu_{2}$ leads to the shear modulus $\mu$. The first Piola-Kirchhoff stress is
$P=\mu_{1}F+\mu_{2}(I_{1}F-F\cdot C).$
Gent model. The Gent model [82] captures material elasticity using logarithm functions. The strain energy density function is
$\Psi(I_{1})=-\frac{\mu J_{m}}{2}\ln(1-\frac{I_{1}-2}{J_{m}}),$
where $\mu$ is the shear modulus and $J_{m}$ is a material stiffening parameter that imposes a maximum allowable stretch on the material. The first Piola-Kirchhoff stress is defined as
$P=-\frac{\mu J_{m}}{J_{m}-(I_{1}-2)}.$
The Green-Lagrangian strains, $\epsilon_{ij}=\frac{1}{2}(F_{ki}F_{kj}-\delta_{ij})$, from finite element simulations were supplied to PINNs as reference data. All FEA analyses were carried out using FEniCS [75].
@@ 5.4 PINNs for heterogeneous material properties identification
As an extension to prior work on utilizing PINNs to determine the macroscopic constitutive parameters in linear and hyperelastic materials [48, 49], we demonstrate a proof-of-concept study to identify the full-field elastic properties across the 2D geometric domain. Our work aims to facilitate future studies to better understand the interconnection between material microscopic properties and its macroscopic mechanical functions. All PINNs models were constructed using the DeepXDE library [85], and the code will be made publicly available after publication in the GitHub repository https://github.com/lu-group/pinn-heterogeneous-material.
@@ 5.4.2 Loss functions
Appendix A: Additional results on method comparison
We present inverse FEA results for the GRF example described in Section 2.1. The optimization process was facilitated using the adjoint method. Specifically, we set to identify a heterogeneous elasticity map that minimizes the $L^{2}$ relative errors of strain under the constraint that $1\leq E\leq 5$. The maximum optimization iteration was set to 100, and the acceptable tolerance of the $L^{2}$ relative error for strains was set to $10^{-3}$. The convergence histories of strain and $E$ in the first 10 iterations are shown in Fig. 10A.
We found that the initialization of the elasticity map, $E_{\text{init}}(x,y)$, has a significant influence on the estimation of the elasticity map. When $E_{\text{init}}=$ 3, 4, and 5, the $L^{2}$ relative errors of strain were consistently around 20% throughout the optimization process. On the other hand, with $E_{\text{init}}(x,y)=$ 1, 2, and randomly distributed, the $L^{2}$ relative errors of strain decreased as the optimization progressed. However, the intermediate estimated elasticity field often resulted in divergence in the FEA, consequently leading to pre-mature termination of the optimization process.
When the initial elastic modulus $E_{\text{init}}(x,y)$ was set to 1, 2, and randomly distributed, a noticeable decrease in the $L^{2}$ relative errors of the strain was evident. However, this decline did not correspond to the $L^{2}$ relative errors of the elasticity map. Surprisingly, an increase in $L^{2}$ relative errors was observed in cases when $E_{\text{init}}(x,y)$ = 2 and randomly distributed, despite a decrease in strain error.
Fig. 10B demonstrates the most accurate estimates, measured by the least $L^{2}$ relative errors, of elasticity maps resulting from the various $E_{\text{init}}(x,y)$. The results indicate that while the inverse FEA roughly uncovered the heterogeneous pattern, the range of elasticity moduli with $E_{\text{init}}(x,y)$ values of 1, 2, 3, 4, and 5 suggests a relatively uniform and close to homogeneous distribution. Notably, when the $E_{\text{init}}(x,y)$ values were randomly distributed, we observed a more reasonable range in elastic moduli. However, the resulting elasticity map estimation did not capture the true pattern of the elasticity map.
Fig. 11 presents the evolutions of the heterogeneous elasticity maps with $E_{\text{init}}(x,y)=$ 1 and randomly distributed. We observed that the optimizer struggled to uncover the true range of the elastic moduli in the case when initializing the elasticity map uniformly to 1. On the contrary, the random initialization of the elasticity map seemed to better account for the wide range of elastic moduli.
Figure 10: Elasticity map convergence trajectories and best estimations. (A) The $L^{2}$ relative errors convergence trajectories of strain and the elasticity map are shown. The $L^{2}$ relative errors in strain consistently remained around 20% with initial values of $E_{\text{init}}(x,y)=$ 3, 4, and 5. The lack of convergence resulted in constant high errors in the elasticity maps, with $L^{2}$ relative errors exceeding 100%. (B) The best estimates of the elasticity map from each elasticity map initialization are presented. The $L^{2}$ relative errors of the estimated elasticity map for $E_{\text{init}}(x,y)$ values of 1, 2, 3, 4, 5, and randomly distributed were 12.93%, 31.84%, 94.67%, 158.87%, 205.18%, and 12.93%.
Figure 11: Elasticity map evolutions. The evolutions of the heterogeneous elasticity maps with $E_{\text{init}}(x,y)=$ 1 and randomly distributed as they provided the best estimates among the numerical experiments. (A) The $L^{2}$ relative errors of the elasticity map at each iteration, when $E_{\text{init}}(x,y)=$ 1, were as follows: 35.75%, 34.80%, 12.93%, 25.49%, 28.45%, and 28.79%. (B) The $L^{2}$ relative errors of the elasticity map at each iteration, with randomly distributed $E_{\text{init}}(x,y)$, were 12.93%, 12.93%, 12.93%, 14.69%, 26.78%, and 29.22%.
Appendix B: Additional results on the selected architectures
Fig. 12 presents the estimated displacements, strain, and elastic modulus distribution of the GRF, brain tissue, and tricuspid valve tissues examples using network architecture IIB. The absolute relative errors between the PINN estimation and FEA ground truth are also provided for comparison. Despite the displacement boundaries being unconstrained in this network architecture, PINN model IIB demonstrated highly accurate estimation of the displacement, strain, and elastic parameter distribution across the three examples with relative errors up to $\mathcal{O}(10^{-2})$ for full-field displacements and strains and $\mathcal{O}(10^{-1})$ for the elastic moduli. These results indicate the consistency and generality of the proposed neural network architecture for estimating the mechanical properties of complex microstructures.
Figure 12: PINN prediction and pointwise error. The FEA ground truth, PINN estimation (denoted with superscript *), and the absolute errors of displacement, strain, and elastic modulus distribution for examples (A), (B), and (C) are shown. The PINN architecture effectively resolves the local features of complex strain and elastic modulus fields. The absolute errors were up to $\mathcal{O}(10^{-2})$ for the displacement and up to $\mathcal{O}(10^{-1})$ for the strain and elasticity fields.
@@ Acknowledgments
This work was supported by the Cora Topolewski Pediatric Valve Center at the Children’s Hospital of Philadelphia, an Additional Ventures Expansion Award, the National Institutes of Health, and the U.S. Department of Energy. WW was initially supported by NHLBI T32 HL007915 and transitioned to NHLBI K25 HL168235. MAJ was supported by NIH R01 HL153166, an Additional Ventures Single Ventricle Grant, and the Topolewski Endowed Chair in Pediatric Cardiology. MD and LL were supported by the U.S. Department of Energy [DE-SC0022953].
@@
The code for this study is available after publication in the GitHub repository https://github.com/lu-group/pinn-heterogeneous-material.